OpenStudy (anonymous):

Determine the number of atoms in a 52.0 gram sample of carbon, how would you set up the conversion?

OpenStudy (anonymous):

@ganeshie8 @satellite73 @chmvijay

OpenStudy (anonymous):

52g/(12g/mol) = # of moles. Convert to atoms using avogadro's constant

OpenStudy (anonymous):

@ECE is this the answer

OpenStudy (anonymous):

Is this the answer

OpenStudy (anonymous):

Avogadro's constant is atoms/mol so yes.

OpenStudy (anonymous):

@chmvijay is this correct I don't quite get it

OpenStudy (chmvijay):

what is one mole of C? how much it weighs?

OpenStudy (anonymous):

idk im so sorry I don't get this

OpenStudy (chmvijay):

one mole of any elements contains =6.023*10^23 of atoms of that element ok

OpenStudy (anonymous):

ok got it

OpenStudy (chmvijay):

so one mole of that element = atomic weight of that element

OpenStudy (anonymous):

let me see what the atomic weight is

OpenStudy (anonymous):

12.01

OpenStudy (anonymous):

ok before we continue what do the lines in an element’s line spectrum represent?

OpenStudy (chmvijay):

ok 12 gram contains 6.023*10^23 atoms of carbon then 52 gram conatins = 6.023*10^23 *52 /12

OpenStudy (chmvijay):

electron transitions of that element it represents

OpenStudy (anonymous):

is that the answer to What do the lines in an element’s line spectrum represent?

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

electron transitions of that element it represents

OpenStudy (anonymous):

ok got it thnx....oh yea what about the other question is this the correct answer "ok 12 gram contains 6.023*10^23 atoms of carbon then 52 gram conatins = 6.023*10^23 *52 /12"

OpenStudy (anonymous):

@chmvijay are you still there

OpenStudy (chmvijay):

yaaa

OpenStudy (chmvijay):

electron transitions of that element it represents

OpenStudy (anonymous):

no im talking about the original question